# A New Upper Bound for Distributed Hypothesis Testing Using the Auxiliary Receiver Approach

### v3

@@ Abstract
This paper employs the add-and-subtract technique of the auxiliary receiver approach to establish a new upper bound for the distributed hypothesis testing problem. This new bound has fewer assumptions than the upper bound proposed by Rahman and Wagner, is at least as tight as the bound by Rahman and Wagner, and can outperform it in certain Gaussian settings. Conceptually speaking, unlike Rahman and Wagner, who view their additional receiver as side information, we view it as an auxiliary receiver and use a different manipulation for single-letterization. [^1] This work was supported by the CUHK Direct Grant 4055193.
@@ I Introduction
This paper concerns the problem of hypothesis testing in a distributed setting. Different variants of the problem are defined in the literature (see [6, 1, 2, 3, 4, 5]). In this paper, we consider the simple two-terminal setup as described in Fig. 1. Let $X$ and $Y$ be two sources with discrete alphabets $\mathcal{X}$ and $\mathcal{Y}$, respectively. Two terminals observe i.i.d. sequences $X^{n}$ and $Y^{n}$ over $n$ time instances. Under the null hypothesis $\mathcal{H}_{0}$, the joint distribution is $P_{XY}$, while under the alternative hypothesis $\mathcal{H}_{1}$, it is $Q_{XY}$. The observed sequence of $X$ is transmitted over an $(n,R)$ channel using an encoding function $f_{n}:\mathcal{X}^{n}\rightarrow\mathcal{M}=\{1,2,\ldots,2^{\lfloor nR\rfloor}\}$. The detector receives the message $m=f_{n}(x^{n})$ and decides between the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ based on the following discriminant rule:
@@ I-A Literature Review
$E_{P_{XY},Q_{XY}}(R)$
$\geq\max_{P_{U|X}:\ I_{P}(U;X)\leq R}\left\{D(P_{X}\|Q_{X})+D(P_{UY}\|Q_{UY})\right\}.$
The use of $P_{U|X}$ implies a Markov chain $U\rightarrow X\rightarrow Y$. The lower bound is not generally tight, as mentioned in [7, Section 3]. The work is then extended and improved by Han [8], and Shimokawa, Han, and Amari (SHA) using the quantization-and-binning technique [9]. Further improvements following this line include [10], [11], [12], [13], and [14].
@@
$E_{P_{XY},Q_{XY}}(R)\leq D(P_{XY}\|Q_{XY}).$
Conversely, at zero rates, Shalaby and Papamacrou [15] demonstrated that transmitting the sequence type is essentially optimal. The authors in [7] established optimality for the special case of testing against independence where $Q_{XY}=P_{X}P_{Y}$. Rahman and Wagner [16] expanded on this by finding the optimal exponent for “testing against conditional independence" setting. Relevant work also includes [17], and [18]. Authors in [16] obtained the following upper bound:
@@
Let $\mathcal{S}(P_{XYZ})$ to be the set of all $P_{U|X}$ such that $P_{UXYZ}=P_{U|X}P_{XYZ}$ satisfies
@@
where $M$ is the transmitted message. In our derivation, we bypass this step and instead utilize a telescopic sum expansion. Consequently, $Z$ does not function as side information but rather as a general auxiliary receiver.
@@
In this paper, we propose a computable upper bound on $E_{P_{XY},Q_{XY}}(R)$ using a systematic manipulation on the auxiliary receivers [21]. The auxiliary receiver approach is a technique for writing new outer bounds for different communication problems. It has been used in [21] to derive new bounds for the broadcast, relay, and interference channels. In this paper, we utilize the “add-and-subtract" technique of the auxiliary receiver approach to derive a new upper bound for the distributed hypothesis testing problem. [^2] The message augmentation technique of the auxiliary receiver approach appears to yield only the bound presented in Proposition 1, which can be proven more readily through a direct method.
@@ I-B Notation
All the logs in this paper are in base $e$. Similarly, all the entropies are measured in nats.
II Upper bound using the add-and-subtract technique
II-A Review of the add-and-subtract technique
@@
    Single-letterize the “difference term" using a telescopic sum expansion. For instance, $I(M;Y^{n})-I(M;Z^{n})$ may be single-letterized as follows:
    $I(M;Y^{n})-I(M;Z^{n})$
    $=\sum_{i}\left(I(M;Y^{i}Z_{i+1}^{n})-I(M;Y^{i-1}Z_{i}^{n})\right).$
@@
    Lastly, reduce the “substituted term" to a single-letter form by following the procedure of an existing upper bound. This is done by mimicking the bound while formally replacing a legitimate receiver with the auxiliary receiver. For example, in the context of a relay channel, consider the cut-set bound and substitute the original receiver $Y$ with the auxiliary receiver $Z$ throughout all the steps of the bound.
This method produces novel upper bounds for relay, broadcast, and interference channels.
A formal proof of applying the add-and-subtract technique to the distributed hypothesis testing problem is presented in Section II-B. In the remainder of this section, we provide an intuitive discussion of the concept.
For the distributed hypothesis testing problem, we consider $D(P_{MY^{n}}\|Q_{MY^{n}})$ as our multi-letter expression (see Theorem 1). After replacing $Y$ with an auxiliary receiver $Z$, we obtain the expression $D(P_{MZ^{n}}\|Q_{MZ^{n}})$ as our “substituted term". Finally, we form and expand the difference term $D(P_{MY^{n}}\|Q_{MY^{n}})-D(P_{MZ^{n}}\|Q_{MZ^{n}})$ via a telescopic sum as follows:
@@
$=\frac{1}{n}\sum_{i=1}^{n}\{D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})$ (3)
$-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})\}$
$=\frac{1}{n}\sum_{i=1}^{n}\left\{D(P_{U_{i}Y_{i}}\|Q_{U_{i}Y_{i}})-D(P_{U_{i}Z_{i}}\|Q_{U_{i}Z_{i}})\right\}$
@@
$\sup_{P_{U|X},Q_{U|X}}D(P_{UY}\|Q_{UY})-D(P_{UZ}\|Q_{UZ}),$
where $P_{U|X}$ and $Q_{U|X}$ satisfy $I_{P}(U;X|Y)\leq R$, $I_{P}(U;X|Z)\leq R$, $I_{Q}(U;X|Y)\leq R$ and $I_{Q}(U;X|Z)\leq R$. We further simplify this expression by dropping the constraints $I_{Q}(U;X|Y)\leq R$ and $I_{Q}(U;X|Z)\leq R$, and by bounding the objective function from above. This allows us to obtain a simpler expression for comparison with previously known bounds. In Section II-B, we define $G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$ to as an upper bound on the difference term. This completes Step 4 of the above procedure and yields:
@@
$=\underbrace{\frac{1}{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D(P_{MY^{n}}\|Q_{MY^{n}})-\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}}_{\text{difference term}}$
$+\underbrace{\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}_{\text{substituted term}}$
$\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+\underbrace{\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}_{\text{substituted term}}$
@@
However, it is not clear if using the centralized upper bound is the best choice.
@@
$\Delta(P_{XYZ},P_{XYJ})=\max_{P_{UV|XY}}I(U;Z|V)-I(U;J|V).$
The term $\Delta(P_{XYZ},P_{XYJ})$ serves as an upper bound on the “difference term", and can be obtained through a telescopic sum expansion. In the case of the key agreement problem, we can complete Step 5. Notably, if we adopt the simple upper bound $S(X;Y\|J)\leq I(X;Y|J)$, we arrive at the following upper bound for $S(X;Y|Z)$, which is currently the best known upper bound for this problem [24]:
@@
II-B The bound obtained using the add-and-subtract technique
@@
Assume that $P_{XYZ}$ and $Q_{XYZ}$, satisfy $D(P_{XY}\|Q_{XY})<\infty$ and $D(P_{XZ}\|Q_{XZ})<\infty$. Then, we have
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+E_{P_{XZ},Q_{XZ}}(R)$
@@
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{n}D(P_{MY^{n}}\|Q_{MY^{n}})-\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}$
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D(P_{MY^{n}}\|Q_{MY^{n}})-D(P_{MZ^{n}}\|Q_{MZ^{n}})}$ (5)
$=\sum_{i=1}^{n}\{D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})$
$-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})\}$ (6)
@@
Let $T\in\{1,2,\cdots,n\}$ be a time-sharing random variable, independent of all previously defined variables and set
@@
$\sum_{i=1}^{n}\{D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})$
$-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})\}$
$=\sum_{i=1}^{n}D(P_{U_{i}Y_{i}}\|Q_{U_{i}Y_{i}})-D(P_{U_{i}Z_{i}}\|Q_{U_{i}Z_{i}})$
$=\sum_{i=1}^{n}\underbrace{D(P_{U_{i}}\|Q_{U_{i}})}_{\text{cancelled}}+D(P_{Y_{i}|U_{i}}\|Q_{Y_{i}|U_{i}}|P_{U_{i}})$
$-\underbrace{D(P_{U_{i}}\|Q_{U_{i}})}_{\text{cancelled}}-D(P_{Z_{i}|U_{i}}\|Q_{Z_{i}|U_{i}}|P_{U_{i}}).$
@@
$\leq n\cdot\sum_{u}P_{U}(u)\max_{\hat{Q}_{X}}\left[D(P_{Y|u}\|\hat{Q}_{Y})-D(P_{Z|u}\|\hat{Q}_{Z})\right]$
where the maximum is over $\hat{Q}_{X}\ll Q_{X}$ (observe that if $Q_{X}(x)=0$, then $Q_{X|u}(x)=0$).
@@
where (9) follows from the i.i.d assumption, and (10) follows from $I(Y^{i-1};X_{i}|Z^{n}X^{i-1}M)=0$ as
@@
where (12) follows from by the data processing inequality, and (13) follows from the definition of $\hat{Q}_{XYZ}$. ∎
We remark that, unlike the left-hand side, the right-hand side in (11) does not depend on $Q_{XZ}$, so one may assume $Q_{XZ}=P_{XZ}$ if using Theorem 5.

### v4

@@ Abstract
This paper employs the add-and-subtract technique of the auxiliary receiver approach to establish a new upper bound for the distributed hypothesis testing problem of distinguishing between the two hypotheses $P_{XY}$ and $Q_{XY}$ using one-way communication. The new bound has fewer assumptions than the upper bound proposed by Rahman and Wagner, is at least as tight as the bound by Rahman and Wagner, and outperforms it in certain discrete and Gaussian settings. Conceptually speaking, unlike Rahman and Wagner, who view their additional receiver as side information, we view it as an auxiliary receiver and use a different manipulation for single-letterization. The new bound also implies that the minimum rate $R$ needed to achieve the centralized bound is at least the conditional entropy $H_{P}(X|Y)$ when $P_{XY}$ and $Q_{XY}$ are non-degenerate distributions.
Index Terms: Auxiliary receivers, distributed hypothesis testing, error exponent.
@@ I Introduction
[…]
@@ I-A Literature Review
$E_{P_{XY},Q_{XY}}(R)\geq\max_{P_{U|X}:\ I_{P}(U;X)\leq R}\left[D(P_{X}\|Q_{X})+D(P_{UY}\|Q_{UY})\right].$
The use of $P_{U|X}$ implies a Markov chain $U\rightarrow X\rightarrow Y$. The lower bound is not generally tight, as mentioned in [1, Section 3]. The work is then extended and improved by Han [2], and Shimokawa, Han, and Amari (SHA) using the quantization-and-binning technique [3]. Further improvements following this line include [4], [5], [6], [7], and [8].
@@
$E_{P_{XY},Q_{XY}}(R)\leq D(P_{XY}\|Q_{XY}).$ (1)
The centralized bound is tight when $R\geq D(P_{XY}\|Q_{XY})+H_{P}(X|Y)$, see [3]. One of the results of this paper is that if the centralized bound is tight, we must have $R\geq H_{P}(X|Y)$. Moreover, in some cases, we can provide better bounds than $H_{P}(X|Y)$.
In addition to characterizing the optimal exponent at high rates when $R\geq D(P_{XY}|Q_{XY})+H_{P}(X|Y)$, the optimal exponent is also known in some other special settings. The authors in [1] found the optimal exponent for the special case of testing against independence where $Q_{XY}=P_{X}P_{Y}$. Rahman and Wagner [9] expanded on this by finding the optimal exponent for the "testing against conditional independence" setting. At zero rate, when we allow for one bit of communication, Han [2] computed the optimal exponent.
In this paper, we are mainly concerned with upper bounds on the exponent $E_{P_{XY},Q_{XY}}(R)$. Authors in [9] obtained the following upper bound:
@@
Let $\mathcal{S}(P_{XYZ})$ be the set of all $P_{U|X}$ such that $P_{UXYZ}=P_{U|X}P_{XYZ}$ satisfies
@@
where $M$ is the transmitted message. In other words, receiver $Y$ is replaced with $Z^{\prime}=(Y,Z)$. In our derivation, we bypass this step and instead utilize a telescopic sum expansion. Consequently, $Z$ does not function as side information but rather as a general auxiliary receiver.
@@
In this paper, we propose a computable upper bound on $E_{P_{XY},Q_{XY}}(R)$ using a systematic manipulation on the auxiliary receivers [12]. The auxiliary receiver approach is a technique for writing new outer bounds for different communication problems. It has been used in [12] to derive new bounds for the broadcast, relay, and interference channels. In this paper, we utilize the "add-and-subtract" technique of the auxiliary receiver approach to derive a new upper bound for the distributed hypothesis testing problem. [^2] The message augmentation technique of the auxiliary receiver approach appears to yield only the bound presented in Proposition 1, which can be proven more readily through a direct method.
Finally, we note that the literature studies different variants and extensions of the problem considered here [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 5, 33, 34, 35]), and our technique may be applied to those settings as well.
@@ I-B Notation
All the logs in this paper are in base $e$, unless stated otherwise. Similarly, all the entropies are measured in nats.
II Upper bound using the auxiliary receiver approach
II-A The auxiliary receiver approach
Here, we review the add-and-subtract idea of the auxiliary receiver approach [12]. Let us view $E(P_{XY},Q_{XY},R)$ as a function of $P_{XY},Q_{XY},R$. The domain of this function is the set of all triples $(P_{XY},Q_{XY},R)$. To study the curvature of this function, we can take another point $(P_{X^{\prime}Y^{\prime}},Q_{X^{\prime}Y^{\prime}},R^{\prime})$ in the domain, and study the difference $E(P_{XY},Q_{XY},R)-E(P_{X^{\prime}Y^{\prime}},Q_{X^{\prime}Y^{\prime}},R^{\prime})$. Figure 2 visualizes this. In this work, we use the second point $(P_{X^{\prime}Y^{\prime}},Q_{X^{\prime}Y^{\prime}},R^{\prime})$ as
$(P_{XZ},Q_{XZ},R)$
so that $X^{\prime}=X,Y^{\prime}=Z,R^{\prime}=R$ for some "auxiliary receiver $Z$". The difference $E(P_{XY},Q_{XY},R)-E(P_{XZ},Q_{XZ},R)$ quantifies the curvature of the plot of the function. We would like to find an upper bound $G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$ on the gap
$E(P_{XY},Q_{XY},R)-E(P_{XZ},Q_{XZ},R)\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R).$
Finding an upper bound on this gap leads to an upper bound on $E(P_{XY},Q_{XY},R)$ via an add-and-subtract manipulation as follows:
$E(P_{XY},Q_{XY},R)$
$=E(P_{XY},Q_{XY},R)-E(P_{XZ},Q_{XZ},R)+E(P_{XZ},Q_{XZ},R)$
$\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+E(P_{XZ},Q_{XZ},R).$
The auxiliary receiver $Z$ is chosen such that either the value of $E(P_{XZ},Q_{XZ},R)$ is known, or an upper bound on it is available. The special case where $Y$ is a function of $Z$ corresponds to a Genie-aided proof. The "side information" approach of Rahman and Wagner [9] is an example of such a choice (the random variable $Z$ in the bound of Rahman and Wagner corresponds to the auxiliary receiver $Z^{\prime}=(Y,Z)$). However, the auxiliary receiver $Z$ does not need to be an enhanced version of $Y$, and this is our main idea to obtain new upper bounds.
$E$$E(P_{XY},Q_{XY},R)$$(P_{XY},Q_{XY},R)$$E(P_{X^{\prime}Y^{\prime}},Q_{X^{\prime}Y^{\prime}},R^{\prime})$$(P_{X^{\prime}Y^{\prime}},Q_{X^{\prime}Y^{\prime}},R^{\prime})$$E(P_{XY},Q_{XY},R)-E(P_{X^{\prime}Y^{\prime}},Q_{X^{\prime}Y^{\prime}},R^{\prime})$
Figure 2:
II-B Review of the add-and-subtract technique
@@
    Single-letterize the "difference term" using a telescopic sum expansion. For instance, $I(M;Y^{n})-I(M;Z^{n})$ may be single-letterized as follows:
    $I(M;Y^{n})-I(M;Z^{n})=\sum_{i}\left[I(M;Y^{i}Z_{i+1}^{n})-I(M;Y^{i-1}Z_{i}^{n})\right].$
@@
    Lastly, reduce the "substituted term" to a single-letter form by following the procedure of an existing upper bound. This is done by mimicking the bound while formally replacing a legitimate receiver with the auxiliary receiver. For example, in the context of a relay channel, consider the cut-set bound and substitute the original receiver $Y$ with the auxiliary receiver $Z$ throughout all the steps of the bound.
This method produces novel upper bounds for relay, broadcast, and interference channels [12].
A formal proof of applying the add-and-subtract technique to the distributed hypothesis testing problem is presented in Section II-C. In the remainder of this section, we provide an intuitive discussion of the concept.
For the distributed hypothesis testing problem, we consider $D(P_{MY^{n}}\|Q_{MY^{n}})$ as our multi-letter expression (see Theorem 1). After replacing $Y$ with an auxiliary receiver $Z$, we obtain the expression $D(P_{MZ^{n}}\|Q_{MZ^{n}})$ as our "substituted term". Finally, we form and expand the difference term $D(P_{MY^{n}}\|Q_{MY^{n}})-D(P_{MZ^{n}}\|Q_{MZ^{n}})$ via a telescopic sum as follows:
@@
$=\frac{1}{n}\sum_{i=1}^{n}\left[D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})\right]$
$=\frac{1}{n}\sum_{i=1}^{n}\left[D(P_{U_{i}Y_{i}}\|Q_{U_{i}Y_{i}})-D(P_{U_{i}Z_{i}}\|Q_{U_{i}Z_{i}})\right]$
@@
$\sup_{P_{U|X},Q_{U|X}}[D(P_{UY}\|Q_{UY})-D(P_{UZ}\|Q_{UZ})]$
where $P_{U|X}$ and $Q_{U|X}$ satisfy $I_{P}(U;X|Y)\leq R$, $I_{P}(U;X|Z)\leq R$, $I_{Q}(U;X|Y)\leq R$ and $I_{Q}(U;X|Z)\leq R$. We further simplify this expression by dropping the constraints $I_{Q}(U;X|Y)\leq R$ and $I_{Q}(U;X|Z)\leq R$, and by bounding the objective function from above. This allows us to obtain a simpler expression for comparison with previously known bounds. In Section II-C, we define $G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$ as an upper bound on the difference term. This completes Step 4 of the above procedure and yields:
@@
$=\underbrace{\frac{1}{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D(P_{MY^{n}}\|Q_{MY^{n}})-\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}}_{\text{difference term}}+\underbrace{\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}_{\text{substituted term}}$
$\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+\underbrace{\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}_{\text{substituted term}}.$
@@
However, it is unclear whether using the centralized upper bound is the best choice.
@@
$\Delta(P_{XYZ},P_{XYJ})=\max_{P_{UV|XY}}[I(U;Z|V)-I(U;J|V)].$
The term $\Delta(P_{XYZ},P_{XYJ})$ serves as an upper bound on the "difference term", and can be obtained through a telescopic sum expansion. In the case of the key agreement problem, we can complete Step 5. Notably, if we adopt the simple upper bound $S(X;Y\|J)\leq I(X;Y|J)$, we arrive at the following upper bound for $S(X;Y|Z)$, which is currently the best known upper bound for this problem [37]:
@@
II-C The bound obtained using the add-and-subtract technique
@@
Assume that $P_{XYZ}$ and $Q_{XYZ}$ satisfy $D(P_{XY}\|Q_{XY})<\infty$ and $D(P_{XZ}\|Q_{XZ})<\infty$. Then, we have
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+E_{P_{XZ},Q_{XZ}}(R).$
@@
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{n}D(P_{MY^{n}}\|Q_{MY^{n}})-\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D(P_{MY^{n}}\|Q_{MY^{n}})-D(P_{MZ^{n}}\|Q_{MZ^{n}})}$
$=\sum_{i=1}^{n}[D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})]$ (5)
@@
Let $T\in\{1,2,\cdots,n\}$ be a time-sharing random variable, independent of all previously defined variables, and set
@@
$\sum_{i=1}^{n}[D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})]$
$=\sum_{i=1}^{n}[D(P_{U_{i}Y_{i}}\|Q_{U_{i}Y_{i}})-D(P_{U_{i}Z_{i}}\|Q_{U_{i}Z_{i}})]$
$=\sum_{i=1}^{n}[\underbrace{D(P_{U_{i}}\|Q_{U_{i}})}_{\text{cancelled}}+D(P_{Y_{i}|U_{i}}\|Q_{Y_{i}|U_{i}}|P_{U_{i}})-\underbrace{D(P_{U_{i}}\|Q_{U_{i}})}_{\text{cancelled}}-D(P_{Z_{i}|U_{i}}\|Q_{Z_{i}|U_{i}}|P_{U_{i}})].$
@@
$\leq n\cdot\sum_{u}P_{U}(u)\max_{\hat{Q}_{X}\ll Q_{X}}\left[D(P_{Y|u}\|\hat{Q}_{Y})-D(P_{Z|u}\|\hat{Q}_{Z})\right]$
where the maximum is over $\hat{Q}_{X}\ll Q_{X}$ (observe that if $Q_{X}(x)=0$, then $Q_{X|u}(x)=0$) and $\hat{Q}_{Y}$, $\hat{Q}_{Z}$ are given by Definition 3.
@@
where (8) follows from the i.i.d. assumption, and (9) follows from $I(Y^{i-1};X_{i}|Z^{n}X^{i-1}M)=0$ as
@@
where (11) follows from the data processing inequality, and (12) follows from the definition of $\hat{Q}_{XYZ}$. ∎
Remark 3.
Please note that, unlike the left-hand side, the right-hand side in (10) does not depend on $Q_{XZ}$, so one may assume $Q_{XZ}=P_{XZ}$ when using Theorem 5.